This is only one problem. All parts please.

Problem 8-6
Expected returns

Stocks A and B have the following probability distributions of expected future returns:

Probability A B
0.1 -7% -38%
0.2 6 0
0.4 14 18
0.2 24 25
0.1 28 47
  1. Calculate the expected rate of return, rB, for Stock B (rA = 13.70%.) Do not round intermediate calculations. Round your answer to two decimal places.
    %

  2. Calculate the standard deviation of expected returns, σA, for Stock A (σB = 21.17%.) Do not round intermediate calculations. Round your answer to two decimal places.
    %

  3. Now calculate the coefficient of variation for Stock B. Round your answer to two decimal places.

  4. Is it possible that most investors might regard Stock B as being less risky than Stock A?  

    1. If Stock B is more highly correlated with the market than A, then it might have the same beta as Stock A, and hence be just as risky in a portfolio sense.
    2. If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.
    3. If Stock B is less highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be more risky in a portfolio sense.
    4. If Stock B is more highly correlated with the market than A, then it might have a higher beta than Stock A, and hence be less risky in a portfolio sense.
    5. If Stock B is more highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.
Answer #1

Expected returns=Sum(probability*returns)

Standard deviation=Sqrt(Sum(probability*(returns-expected returns)^2))

Coefficient of variation=Standard deviation/expected returns

1.
=0.1*(-38%)+0.2*0%+0.4*18%+0.2*25%+0.1*47%
=13.100%

2.
=sqrt(0.1*(-7%-13.70%)^2+0.2*(6%-13.70%)^2+0.4*(14%-13.70%)^2+0.2*(24%-13.70%)^2+0.1*(28%-13.70%)^2)
=9.82%

3.
=21.17%/13.100%
=1.616030534

4.
If Stock B is less highly correlated with the market than A, then it might have a lower beta than Stock A, and hence be less risky in a portfolio sense.